OpenStudy (anonymous):

Dustin's starting salary is $40,000. His salary is expected to increase by 4% each year. Which expression can be used to calculate Dustin's salary n years from now? To the nearest dollar, what is Dustin's expected salary in 5 years? A) 40000(1.04)(n - 1) ; $46,794.00 B) 40000(1.04)n ; $48,666.00 C) 40000(1.04)(n + 1) ; $50,613.00 D) 40000(1.04)(n - 2) ; $44,795.00

OpenStudy (zyi6):

Do you understand what the question is asking at all?

OpenStudy (zyi6):

well, this is the explanation on how to solve it. First, we have to set up our exponential function: f(n) = 40,000(1.04)^n Next, we substitute 5 in for the exponent. f(n) = 40,000(1.04)^5 Then, solve. f(n) = 40,000(1.04)^5 f(n) = 40,000(1.2166529024) f(n) = 48666.116096 This can be rounded to say f(n) = 48,666.00 So, B is your answer. Do you understand how I did that?